OpenStudy (anonymous):

What is the sample standard deviation? Data Values Frequency 3.0 - 6.9 5 7.0 - 10.9 4 11.0 - 14.9 8 15.0 - 18.9 6 19.0 - 22.9 5

OpenStudy (amistre64):

\[variance=\frac{\sum(x-\bar x)^2}{n-1}\] sd = \(\sqrt{variance}\)

OpenStudy (amistre64):

find the mean

OpenStudy (amistre64):

\[mean=\frac{\sum\ (5P(x_1)+4P(x_2)+8P(x_3)+6P(x_4)+5P(x_5))}{5+4+8+6+5}\]

OpenStudy (amistre64):

where any given P(x) is the average of the class interval i believe

OpenStudy (amistre64):

(5*4.95+4*35.8+8*12.95+6*16.95+5*20.95) = 478 478/28 =abt 17.0714

OpenStudy (amistre64):

(5-17.0714)^2+(4-17.0714)^2+(8-17.0714)^2+(6-17.0714)^2+(5-17.0714)^2 = 667.16509/27 = 24.71 sd = sqrt(24.71) = 4.971 if i recalled it correctly